Findall values of that satisfy the equation: 5x4 - 5x3 - 1x² = 0 Solve:
Answer #1

salh soch-5x²_1x² = 0 5x²x² - 5x² x - 1x²=0 x²[5 x ² - 50-1] =0 x=0 ie, x, =0, x=0 Now 5x²5x-1=0 x=0,0 we get two value of a